Rugged Sports Enterprises LP is organized as a limited partnership consisting of two individual partners: Hockey LP and Football LP. Both partners separately operate a minor league hockey team and a semipro football team. Hockey LP Football LP Total Balance at 6/30/2016 $ 338,000 $ 1,385,800 $ 1,723,800 Annual net income 46,554 954,357 1,000,911 Cash distribution — (85,000 ) (85,000 ) Balance at 6/30/2017 $ 384,554 $ 2,255,157 $ 2,639,711 Compute partner return on equity for each limited partnership (and the total) for the year ended June 30, 2017, using the following selected data on partner capital balances from Rugged Sports Enterprises LP.
Answer #1

1.

Return on equity
Numerator / Denominator = Return on equity
Net income / Average equity = Return on equity
Rugged sports enterprises LLP 1000911 / 2181755.5 = 45.88%

Average equity = (1723800+2639711)/2 = 2181755.5

2.

Partner Return on equity
Numerator / Denominator = Return on equity
Net income / Average equity = Return on equity
Hockey LP 46554 / 361277 = 12.89%
Football LP 954357 / 1820478.5 = 52.42%

Hockey LP average equity = (338000+384554)/2 = 361277

Football LP average equity = (1385800 +2255157 )/2 = 1820478.5